OpenStudy (anonymous):

Identifying domain help? Using interval notation, identify the domain for the function: f(x) = (sqrt)x+4 (A) [4, infinity) (B) (-infinity, 4] (C) [-4, infinity) (D) (-infinity, -4] Leaning towards A or C for starters because the bracket should be on the left side when listing the domain, right?

OpenStudy (anonymous):

i wanna help her ermm

ganeshie8 (ganeshie8):

thats a good start!

OpenStudy (anonymous):

i need to know her guess first :O

OpenStudy (anonymous):

I'm not sure how the four can become a negative in this equation so to take a stab at the wind I'd say A @Marki

OpenStudy (anonymous):

wrong

ganeshie8 (ganeshie8):

\[f(x) = \sqrt{x+4}\] assuming everything is real, you want the stuff inside radical to be not less than 0 because squareroot is not defined for negative numbers

OpenStudy (anonymous):

Wouldn't that leave C as my only option?

ganeshie8 (ganeshie8):

\[\large x+4 \ge 0\] solve \(x\)

OpenStudy (anonymous):

x => -4?

OpenStudy (anonymous):

ahh I see now...

ganeshie8 (ganeshie8):

yes, in interval notation it would be \[\large \left[-4, \infty \right)\]

OpenStudy (anonymous):

thank you so much, just got one more for now